OpenStudy (anonymous):

Solve x2 + 10x + 7 = 0 by completing the square. Which equation is used in the process? x + 10)2 = 93 (x + 5)2 = 32 (x + 5)2 = 18

OpenStudy (anonymous):

\[x^2+10x=-7\]\[x^2+10x+25=-7+25\]\[(x+5)^2=18\]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

You are surely welcome! But do you understand what I did?

OpenStudy (anonymous):

Somewhat where did the 25 come from?

OpenStudy (anonymous):

to complete the square you add \[(10/2)^2\] or in the form ax^2+bx+c (b/2)^2=c

OpenStudy (anonymous):

Do you understand?